OpenStudy (nightmarenight):

Help? Will fan and give medal for right answer!! If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4)

OpenStudy (ybarrap):

Just in case you still need help. $$ f(x)=kx^2\\ f(5)=k5^2=75\\ k=\cfrac{75}{25}=3\\ f(x)=3x^2\\ f(4)=3(4)^2=3\times16=42 $$

OpenStudy (nightmarenight):

the answer is 42? that's not an option o: @ybarrap

OpenStudy (ybarrap):

3*16=48 not 42